OpenStudy (maheshmeghwal9):

IF \[\color{red}{|z|=1.}\]&\[\color{red}{z \ne \pm 1.}\]then all the values of \[\LARGE{\frac{z}{1-z^2}.}\]lie on ?

OpenStudy (maheshmeghwal9):

z = a complex number.

OpenStudy (maheshmeghwal9):

Its answer is imaginary axis.

OpenStudy (maheshmeghwal9):

bt hw to gt it? I don't know:(

OpenStudy (anonymous):

\[z=re^{i \theta} \\ r=1 \\ z=e^{i \theta}\\ \frac{z}{1-z^2}=\frac{e^{i \theta}}{1-e^{2i \theta}}\]

OpenStudy (maheshmeghwal9):

\[\frac{e^{i \theta }}{1-e^{2i \theta}}\]\[\implies \frac{e^{i \theta}}{(1+e^{i \theta })(1-e^{i \theta})}.\]\[\implies ?\]

OpenStudy (maheshmeghwal9):

sorry i didn't gt:(

OpenStudy (anonymous):

\[e^{i \theta}-e^{-i \theta}=2i \sin \theta\]

OpenStudy (maheshmeghwal9):

no no that nt actually i didn't gt how come u gt the following: -\[e^{i \theta}-e^{- i \theta}\]

OpenStudy (anonymous):

oh sorry Multiplying Fraction by e^(-i theta) \[\frac{e^{i \theta}}{1-e^{2i \theta}}=\frac{e^{i \theta}e^{-i \theta}}{e^{-i \theta}-e^{-i \theta}e^{2i \theta}}\]

OpenStudy (maheshmeghwal9):

so i get \[- \frac{1}{2\space i \space im (z)}.\]now wt?

OpenStudy (anonymous):

well thats your answer \[-\frac{1}{2i \sin \theta}=-\frac{i}{2i^2 \sin \theta}=\frac{i}{2\sin \theta}=ib \\ b \in \mathbb{R} \\ \because \\ -1 \le \sin \theta \le1 \\ then \\ -\infty \le\frac{1}{\sin \theta} \le \infty \]

OpenStudy (maheshmeghwal9):

so (z) can lie on any point in whole imaginary axis, can't it?

OpenStudy (anonymous):

not z but z/1-z^2

OpenStudy (maheshmeghwal9):

oops sorry my bad. k btw thanx a lot i gt it:)

OpenStudy (anonymous):

ur welcome dear